(41) find the equation of the plane passing through the points A(3,-2,0), B(2,0,3), and C(1,-1,1) in

(i) vector form

(ii) parametric form

(iii) cartesian form
Answer #1

C 1 B - (2,0,3) (1,-1, A = (3.-2,0) & Vector Equation - 0 1 1-2 2 vor bra – 11 +27 +3k e - à -27 + ŷ tk now (5-2) x (e-a = 16 Parameth toon x + y ) + 2 k = (37_29) + M(-) + 2+ 3x) + + (-21+ jatk x= 3-4-at y = -2+2 4+ t 0 +34 tt